### v2

# Robust Point Matching with Distance Profiles

Abstract
While matching procedures based on pairwise distances are conceptually appealing and thus favored in practice, theoretical guarantees for such procedures are rarely found in the literature. We propose and analyze matching procedures based on distance profiles that are easily implementable in practice, showing these procedures are robust to outliers and noise. We demonstrate the performance of the proposed method using a real data example and provide simulation studies to complement the theoretical findings.
1 Introduction
Matching is one of the most fundamental tasks of modern data science, which can be found in numerous applications. For instance, the goal of shape correspondence problems [Van Kaick et al., 2011] in computer graphics is to match corresponding parts of different geometric shapes, often obtained by 3D scanners or imaging equipment used for Magnetic Resonance Imaging (MRI) or Computed Tomography (CT). Graph matching aims to find corresponding nodes between two or more graphs representing structured data, such as neuronal connectivity in human brains [Sporns et al., 2005] or social networks [Borgatti et al., 2009].
A formal mathematical framework for matching [Mémoli, 2011] conceptualizes objects as metric measure spaces, which are essentially probability measures defined on suitable metric spaces. For instance, a smooth 3D surface can be viewed as a uniform measure supported on the surface in the Euclidean space $\mathbb{R}^{3}$. In practice, of course, objects are not fully available and thus approximated by suitable point clouds, say, $\{X_{1},\ldots,X_{n}\}$ and $\{Y_{1},\ldots,Y_{m}\}$ on $\mathbb{R}^{d}$, respectively, and the practical goal is to match these point clouds. The most standard approach is quadratic matching which formulates and solves a quadratic program based on pairwise distances $\{\|X_{i}-X_{\ell}\|_{2}\}_{i,\ell=1}^{n}$ and $\{\|Y_{j}-Y_{\ell}\|_{2}\}_{j,\ell=1}^{m}$. When $n=m$, quadratic matching finds a one-to-one correspondence that aligns the pairwise distances as closely as possible, which is formulated as a Quadratic Assignment Problem (QAP) [Koopmans and Beckmann, 1957, Loiola et al., 2007]. When $n\neq m$, the notion of correspondence is relaxed to soft assignments or couplings, leading to the formulation of the Gromov-Wasserstein (GW) distance [Mémoli, 2011], which has been increasingly used in applications, such as word alignment [Alvarez-Melis and Jaakkola, 2018], statistical inference [Brécheteau, 2019, Weitkamp et al., 2022], and generative modeling [Bunne et al., 2019, Hur et al., 2024]. On the practical side, however, quadratic matching requires solving a quadratic program that is NP-hard in general, and thus needs to rely on several methods, such as convex relaxation [Aflalo et al., 2015], semidefinite relaxation [Zhao et al., 1998, Kezurer et al., 2015], spectral relaxation [Leordeanu and Hebert, 2005], or entropic regularization [Rangarajan et al., 1999, Solomon et al., 2016, Peyré et al., 2016].
Matching procedures based on the pairwise distances $\{\|X_{i}-X_{\ell}\|_{2}\}_{i,\ell=1}^{n}$ and $\{\|Y_{j}-Y_{\ell}\|_{2}\}_{j,\ell=1}^{m}$ are invariant to rigid transformations, which is desirable in practice. For instance, in imaging problems, the images of the objects of interest are often obtained from different angles and thus need to be matched up to rigid transformations. However, such procedures are difficult to analyze theoretically and the literature has been mostly establishing theoretical guarantees for other matching procedures using $\|X_{i}-Y_{j}\|_{2}$’s, which aim to match the points based on the proximity. [Collier and Dalalyan, 2016, Kunisky and Niles-Weed, 2022, Wang et al., 2022] establish theoretical results on recovering the ground truth matching for the procedures based on $\|X_{i}-Y_{j}\|_{2}$’s under some noise and [Galstyan et al., 2022, Minasyan et al., 2023, Galstyan and Minasyan, 2023] show matching guarantees in the presence of outliers. Though Theorem 2 of [Wang et al., 2022] is concerned with a procedure using the pairwise distances $\{\|X_{i}-X_{\ell}\|_{2}\}_{i,\ell=1}^{n}$ and $\{\|Y_{j}-Y_{\ell}\|_{2}\}_{j,\ell=1}^{m}$, it requires solving a nontrivial minimization over permutations, which is computationally challenging, and the result does not consider the presence of outliers.
Motivated by the above challenges, this paper theoretically establishes outlier robustness and noise stability for a matching procedure based on the pairwise distances $\{\|X_{i}-X_{\ell}\|_{2}\}_{i,\ell=1}^{n}$ and $\{\|Y_{j}-Y_{\ell}\|_{2}\}_{j,\ell=1}^{m}$, which is easily implementable in practice. The proposed procedure associates each point with the set of distances from it to other points, namely, $X_{i}$ and $Y_{j}$ are associated with the sets of distances to the other points $\{\|X_{i}-X_{\ell}\|_{2}\}_{\ell=1}^{n}$ and $\{\|Y_{j}-Y_{\ell}\|_{2}\}_{\ell=1}^{m}$, which we call the distance profiles of $X_{i}$ and $Y_{j}$, respectively. The matching procedure is to find for each $X_{i}$, the point $Y_{j}$ having the most similar distance profile to that of $X_{i}$, where the similarity is measured by a suitable quantity between the sets $\{\|X_{i}-X_{\ell}\|_{2}\}_{\ell=1}^{n}$ and $\{\|Y_{j}-Y_{\ell}\|_{2}\}_{\ell=1}^{m}$. The idea of associating a point with suitable invariants based on the distances to other points from it has a long history in the object matching literature, which includes shape contexts [Belongie et al., 2002, Ruggeri and Saupe, 2008], integral invariants [Manay et al., 2006], invariant histograms [Brinkman and Olver, 2012], and the local distribution of distances [Mémoli, 2011]. However, theoretical results on robustness of such methods are rarely found in the literature. In the context of random graph matching, [Ding et al., 2021, Ding et al., 2023] propose and analyze a degree profile matching procedure based on the adjacency matrices $(A_{i\ell})_{i,\ell=1}^{n}$ and $(B_{j\ell})_{j,\ell=1}^{n}$ of certain correlated random graphs. However, since their analysis focuses on random graph models, such as Erdös-Rényi and correlated Wigner models, where the edges are independently generated, $(A_{i\ell})_{i,\ell=1}^{n}$ are mutually independent, which is substantially different from our setting where the distances $\{\|X_{i}-X_{\ell}\|_{2}\}_{i,\ell=1}^{n}$ are not mutually independent.
Our contribution
We show the outlier robustness and noise stability of the matching procedure based on distance profiles, which is easy to implement in practice. Though the main principle of the proposed procedure, namely, matching the points via certain distance-based invariants assigned to them, has already been proposed and rediscovered multiple times in the literature as we mentioned earlier, there has been no theoretical analysis on the outlier robustness and noise stability of this procedure to the best of our knowledge. Under certain probabilistic settings, we prove that the proposed matching procedure is successful with high probability even in the presence of outliers and noise. We demonstrate the performance of the proposed method using a real data example and provide simulation studies to complement the theoretical findings.
Notation
For any $k\in\mathbb{N}$, let $[k]$ denote $\{1,\ldots,k\}$, $\Delta_{k}=\{(p_{1},\ldots,p_{k})\in\mathbb{R}_{+}^{k}:\sum_{i=1}^{k}p_{i}=1\}$ denote the probability simplex, and $\mathcal{S}_{k}$ denote the set of all permutations on $[k]$. For $a,b\in\mathbb{R}$, let $a\vee b=\max\{a,b\}$ and $a\wedge b=\min\{a,b\}$. Let $\mathscr{P}(\mathbb{R}^{d})$ denote the set of all Borel probability measures defined on $\mathbb{R}^{d}$ and for any $z\in\mathbb{R}^{d}$, let $\delta_{z}\in\mathscr{P}(\mathbb{R}^{d})$ denote the Dirac measure at $z$. Lastly, we denote by $W_{1}$ the Wasserstein-1 distance.
1.1 Distance Profile Matching
Algorithm 1 describes the distance profile matching procedure that we propose to study. First, notice that Algorithm 1 is implementable without knowing the points $X_{i}$’s and $Y_{j}$’s as long as the pairwise distances $\{\|X_{i}-X_{\ell}\|_{2}\}_{i,\ell=1}^{n}$ and $\{\|Y_{j}-Y_{\ell}\|_{2}\}_{j,\ell=1}^{m}$ are provided. The discrepancy measure $\mathcal{D}$ quantifies the similarity between the distance profiles. Therefore, for each $X_{i}$, the procedure (1.2) is finding $Y_{j}$ whose distance profile is closest to that of $X_{i}$ under $\mathcal{D}$.
The output $\pi$ of Algorithm 1 represents the matching: $X_{i}$ is matched to $Y_{\pi(i)}$. Note that $\pi$ may not be injective or surjective, meaning that some points may not be matched or some points may be matched to multiple points. [^1] If $n=m$ and one wants to find a one-to-one correspondence, Algorithm 1 can be modified to output a permutation as we discuss in Section 3. The motivation behind this is simple: (1) in the presence of outliers, some points can be ignored or matched arbitrarily, and (2) similar points can be matched to the same point. The underlying objects that the points $X_{i}$’s and $Y_{j}$’s approximate, say, probability measures $\mu$ and $\nu$ defined on $\mathbb{R}^{d}$, respectively, are not identical in practice. Usually, they coincide up to several outlier parts, which we can formally state as $\mu=\lambda\rho+(1-\lambda)\mu_{o}$ and $\nu=\lambda\rho+(1-\lambda)\nu_{o}$ for some $\lambda\in(0,1)$ and $\rho,\mu_{o},\nu_{o}\in\mathscr{P}(\mathbb{R}^{d})$. Here, $\lambda\rho$ represents the common part of $\mu$ and $\nu$ that we want to match, while $(1-\lambda)\mu_{o}$ and $(1-\lambda)\nu_{o}$ are the outlier parts of $\mu$ and $\nu$, respectively. We are mainly interested in matching $X_{i}$’s and $Y_{j}$’s from the common part, while ignoring or arbitrarily matching the points from the outlier parts.
The threshold $\rho$ is an optional parameter that can be used to obtain the subset $I\subset[n]$ containing the indices $i$’s such that the discrepancy $D(i,\pi(i))$ between the distance profiles of $X_{i}$ and its match $Y_{\pi(i)}$ is below $\rho$. Providing a small threshold leads to a more conservative matching that only focuses on the points that are closely matched, which can be used for other downstream tasks. For instance, one can solve the orthogonal Procrustes problem using the matched pairs $\{(X_{i},Y_{\pi(i)})\}_{i\in I}$ to find the optimal rotation and translation aligning them as closely as possible, which theoretically works with just $|I|=d$ pairs in $\mathbb{R}^{d}$ but practically requires more pairs, say, $|I|=O(d)$, due to potential noise or degeneracy. The obtained transformation can then be used as an initialization of the registration procedures, such as the Iterative Closest Point (ICP) algorithm [Besl and McKay, 1992], to find the optimal transformation that aligns the whole point clouds.
In this paper, we mainly use the Wasserstein-1 distance $W_{1}$ as the discrepancy measure so that $D(i,j)$ in (1.2) becomes
$W(i,j):=W_{1}\left(\frac{1}{n}\sum_{\ell=1}^{n}\delta_{\|X_{i}-X_{\ell}\|_{2}},\frac{1}{m}\sum_{\ell=1}^{m}\delta_{\|Y_{j}-Y_{\ell}\|_{2}}\right)\quad\forall(i,j)\in[n]\times[m].$ (1.1)
The Wasserstein-1 distance between two sets of points in $\mathbb{R}$ admits a closed-form expression based on the corresponding cumulative distribution functions, which can be computed efficiently [Peyré and Cuturi, 2019].
Algorithm 1 Distance Profile Matching
0:  Pairwise distances $\{\|X_{i}-X_{\ell}\|_{2}\}_{i,\ell=1}^{n}$ and $\{\|Y_{j}-Y_{\ell}\|_{2}\}_{j,\ell=1}^{m}$.
0:  A well-defined discrepancy measure $\mathcal{D}$ between two sets of points in $\mathbb{R}$.
0:  (Optional) A threshold $\rho>0$. Set $\rho=\infty$ if not provided.
1:  For each pair $(i,j)\in[n]\times[m]$, define
$D(i,j)=\mathcal{D}(\{\|X_{i}-X_{\ell}\|_{2}\}_{\ell=1}^{n},\{\|Y_{j}-Y_{\ell}\|_{2}\}_{\ell=1}^{m}).$
2:  Define a map $\pi\colon[n]\to[m]$ by
$\pi(i)=\operatorname*{arg\,min}_{j\in[m]}D(i,j).$ (1.2)
3:  Define $I\subset[n]$ by $I=\{i\in[n]:D(i,\pi(i))<\rho\}$.
4:  return  $\pi$ and $I$.
1.2 Motivating Application
Before presenting the main theoretical results, we demonstrate how the proposed distance profile matching procedure performs in practice using a real data example. We apply the distance profile matching method to the shape correspondence using CAPOD data [Papadakis, 2014]. Figure 1(a) and Figure 1(d) show the two 3D images representing dolphins with different poses, from which we obtain point clouds $X_{1},\ldots,X_{n}$ and $Y_{1},\ldots,Y_{m}$ as shown in Figure 1(b) and Figure 1(e), respectively, where $n=m=285$ in this example. We want to match the points corresponding to the same part of the dolphins, rather than finding a one-to-one correspondence. As explained earlier, it is reasonable to match similar points to the same point. For instance, there are many points concentrated around the nose—called the rostrum—of the source dolphin, which are expected to be matched to any points concentrated around the rostrum of the target dolphin. Hence, we first divide the dolphins into several components. Here, we use $k$-means clustering with $k=7$, which indeed cluster points based on the body parts. Figure 1(e) shows the target points $Y_{j}$’s colored by $s(j)$’s, where $s(j)\in\{1,\ldots,7\}$ is the index of the component that $Y_{j}$ belongs to. Now, we run Algorithm 1 with $\mathcal{D}=W_{1}$ and obtain the output $\pi\colon[n]\to[m]$. For each $X_{i}$, we verify if the part that it belongs to coincides with the part that its match $Y_{\pi(i)}$ belongs to, which is indexed by $s(\pi(i))$. To this end, Figure 1(b) shows the source points $X_{i}$’s colored by $s(\pi(i))$’s. We can observe that the proposed procedure recovers the corresponding regions mostly correctly. For instance, for $X_{i}$’s corresponding to the rostrum of the source dolphin, their colors $s(\pi(i))$’s (green) indicate that they are matched to the points corresponding to the rostrum of the target dolphin. Similarly, the points corresponding to the head (lime), tail (red), dorsal fin (blue), and peduncle (magenta) are mostly matched correctly, while there are some mismatches, for instance, among the points around the pectoral fin.
Figure 1(c) and Figure 1(f) repeat the same procedure after removing 16 points around the fin of the target dolphin in Figure 1(e) and clustering again with $k=6$. Figure 1(c) shows that the matching result is similar to the previous one. In this case, though the points around the fin of the source dolphin do not have corresponding points in the target dolphin and thus we may ignore matching them, the proposed procedure still matches them to the points around the peduncle (magenta) of the target dolphin, which is a cluster adjacent to the removed fin.
(a) Source object
(b) Source points (color = $s(\pi(i))$)
(c) Source points (color = $s(\pi(i))$)
(d) Target object
(e) Target points (color = $s(j)$)
(f) Target points (color = $s(j)$)
Figure 1: 3D images of dolphins and the corresponding point clouds. (a) and (d) are the source and target 3D images, from which we obtain point clouds $X_{1},\ldots,X_{n}$ and $Y_{1},\ldots,Y_{m}$, with $n=m=285$, as shown in (b) and (e), respectively. We apply $k$-means clustering to both point clouds with $k=7$ to cluster them based on the body parts. (e) shows the target points $Y_{j}$’s colored by $s(j)$’s, where $s(j)\in\{1,\ldots,7\}$ denotes the index of the component that $Y_{j}$ belongs to. (b) shows the source points $X_{i}$’s colored by $s(\pi(i))$’s, where $\pi$ is obtained by applying Algorithm 1 with $\mathcal{D}=W_{1}$. (c) and (f) repeat (b) and (e), respectively, after removing 16 points around the fin of the target dolphin in (e) and clustering again with $k=6$.
2 Outlier Robustness of Distance Profile Matching
This section analyzes the outlier robustness of the distance profile matching procedure described in Algorithm 1. As we have seen in the previous section, Algorithm 1 is designed to match the points corresponding to the same region of the objects, while ignoring or arbitrarily matching the points from the outlier parts. In order to theoretically analyze such outlier robustness, we consider appropriate model assumptions on the underlying objects $\mu,\nu\in\mathscr{P}(\mathbb{R}^{d})$ that capture the nature of the outliers. Motivated by several practical considerations, we model $\mu,\nu$ as follows.
Assumption 1.
Suppose that $\mu$ and $\nu$ are mixture distributions consisting of $t$ and $s$ components, respectively, defined as follows: for some $(p_{1},\ldots,p_{t})\in\Delta_{t}$ and $(q_{1},\ldots,q_{s})\in\Delta_{s}$,
$\mu=\sum_{k=1}^{t}p_{k}N(\theta_{k},\sigma_{k}^{2}I_{d})\quad\text{and}\quad\nu=\sum_{k=1}^{s}q_{k}N(\eta_{k},\tau_{k}^{2}I_{d}),$
For some $K\leq\min\{t,s\}$, we assume that the first $K$ components of $\mu$ and $\nu$ are identical, that is, $p_{k}=q_{k}$, $\theta_{k}=\eta_{k}$, and $\sigma_{k}=\tau_{k}$ for $k\in[K]$.
In other words, Assumption 1 is equivalent to letting $\mu=\lambda\rho+(1-\lambda)\mu_{o}$ and $\nu=\lambda\rho+(1-\lambda)\nu_{o}$, where the common part $\lambda\rho$ is the mixture consisting of the first $K$ components and the outlier parts $(1-\lambda)\mu_{o},(1-\lambda)\nu_{o}$ are the mixtures of the remaining components. The motivation behind Assumption 1 is that the objects $\mu,\nu$ are composed of several regions as we have seen from the example in Section 1.1, where each region is represented as a Gaussian distribution so that the mean corresponds to the center of the region and the variance roughly represents the size of the region. The points corresponding to the same region are to be matched, while we ignore matching the points from the outlier components.
Now, we apply Algorithm 1 to the point clouds $\{X_{1},\ldots,X_{n}\}$ and $\{Y_{1},\ldots,Y_{m}\}$ that approximate the objects $\mu$ and $\nu$, respectively. We take a viewpoint that the points $X_{1},\ldots,X_{n}$ and $Y_{1},\ldots,Y_{m}$ are independent samples from $\mu$ and $\nu$, respectively. For $i\in[n]$ and $j\in[m]$, let $t(i)\in[t]$ and $s(j)\in[s]$ denote the components where $X_{i}$ and $Y_{j}$ are from, respectively. Our goal is to match $X_{i}$ and $Y_{j}$ if $t(i)=s(j)\in[K]$, namely, they are from the same region. Under this probabilistic setting, we show that Algorithm 1 outputs the matching $\pi$ that has the following guarantee: for $X_{i}$ such that $t(i)\in[K]$, we have $s(\pi(i))=t(i)$, namely, its match $Y_{\pi(i)}$ is from the corresponding component. More precisely, we show that the event
$\bigcap_{i\in[n]:t(i)\in[K]}(s(\pi(i))=t(i)),$
which represents the successful matching of the points from the common components, holds with high probability.
There are several factors determining the success of the matching. First, the objects $\mu,\nu$ must be well-configured in that the centers avoid certain symmetries that cannot be distinguished by any procedures based on pairwise distances. For instance, suppose $K=3$ and $\|\theta_{k}-\theta_{k^{\prime}}\|_{2}$ is the same for any $k\neq k^{\prime}$, meaning that all components are equidistant from each other. Then, any matching procedure based on $\|X_{i}-X_{\ell}\|_{2}$’s and $\|Y_{j}-Y_{\ell}\|_{2}$’s, such as the proposed one or the usual quadratic matching, essentially treats all the points as the same point and thus fail to distinguish the components. Therefore, we need a suitable assumption on the configuration of the centers for matching to be well-defined. In addition to proper configurations, the sample sizes $n,m$ must be sufficiently large to approximate $\mu,\nu$ by the samples. Meanwhile, the variances $\sigma_{k}$’s and $\tau_{k}$’s must be sufficiently small to ensure that the mixture components are well-separated from each other. The sample sizes also affect how small the variances should be. In fact, the larger the sample sizes $n,m$ are, the smaller the variances should be to uniformly control the deviation of the samples from their centers. The following theorem formally shows the matching guarantee where the aforementioned conditions on the configurations, sample sizes, and variances are stated as (i), (ii), and (iii), respectively.
Theorem 1.
Suppose $\mu$ and $\nu$ satisfy Assumption 1. Let $X_{1},\ldots,X_{n}$ and $Y_{1},\ldots,Y_{m}$ be independent samples from $\mu$ and $\nu$, respectively. Let $\pi\colon[n]\to[m]$ be the output of Algorithm 1 with $\mathcal{D}=W_{1}$. For $i\in[n]$ and $j\in[m]$, let $t(i)\in[t]$ and $s(j)\in[s]$ denote the components where $X_{i}$ and $Y_{j}$ are from, respectively. Then, for $\delta\in(0,1)$ such that $\delta\leq 12e^{-4(t\vee s)/5}$, we have
$\mathbb{P}\left(\bigcap_{i\in[n]:t(i)\in[K]}(s(\pi(i))=t(i))\right)\geq 1-\delta$ (2.1)
if the following conditions hold.
  • (i)
    Letting
    $\overline{W}(\alpha,\beta):=W_{1}\left(\sum_{k=1}^{t}p_{k}\delta_{\|\theta_{\alpha}-\theta_{k}\|_{2}},\sum_{k=1}^{s}q_{k}\delta_{\|\eta_{\beta}-\eta_{k}\|_{2}}\right)\quad\forall(\alpha,\beta)\in[t]\times[s],$
    suppose
    $\min_{\alpha\neq\beta}\overline{W}(\alpha,\beta)-\max_{\alpha\in[K]}\overline{W}(\alpha,\alpha)=:\Omega>0.$ (2.2)
  • (ii)
    Letting $R=\max_{\alpha,\beta\in[t]}\|\theta_{\alpha}-\theta_{\beta}\|_{2}\vee\max_{\alpha,\beta\in[s]}\|\eta_{\alpha}-\eta_{\beta}\|_{2}$, suppose
    $n\wedge m\geq\frac{1600R^{2}\log(12/\delta)}{\Omega^{2}}.$ (2.3)
  • (iii)
    Letting $\sigma=\max_{k\in[t]}\sigma_{k}$ and $\tau=\max_{k\in[s]}\tau_{k}$, suppose
    $\sigma^{2}\leq\frac{\Omega^{2}}{256(d\vee 8\log(4n^{2}/\delta))}\quad\text{and}\quad\tau^{2}\leq\frac{\Omega^{2}}{256(d\vee 8\log(4m^{2}/\delta))}.$ (2.4)
As discussed earlier, the condition (i) of Theorem 1 ensures that the centers of the mixture components are well-configured so that the matching is well-defined. Particularly, (2.2) ensures that for any $\alpha\in[K]$,
$\operatorname*{arg\,min}_{\beta\in[s]}\overline{W}(\alpha,\beta)=\alpha,$
implying that $\eta_{\alpha}$ is closest to $\theta_{\alpha}$ in terms of the distance profiles $\{\|\theta_{\alpha}-\theta_{k}\|_{2}\}_{k=1}^{t}$ and $\{\|\eta_{\alpha}-\eta_{k}\|_{2}\}_{k=1}^{s}$ properly weighted by $(p_{1},\ldots,p_{t})$ and $(q_{1},\ldots,q_{s})$, respectively, under $W_{1}$. The condition (ii) ensures that the sample sizes $n,m$ are sufficiently large to approximate $\mu,\nu$. The condition (iii) characterizes how small the variances $\sigma_{k}$’s and $\tau_{k}$’s should be given the sample sizes $n,m$ and $\Omega$.
Proof of Theorem 1.
Note that $D(i,j)$ of Algorithm 1 is $W(i,j)$ defined in (1.1) as we let $\mathcal{D}=W_{1}$. For any pair $(i,j)\in[n]\times[m]$, we show that $W(i,j)\approx\overline{W}(t(i),s(j))$ holds. To this end, we define an intermediate quantity $W^{\prime}(i,j)$ as follows: for any $(i,j)\in[n]\times[m]$,
$\begin{split}W^{\prime}(i,j)&:=W_{1}\left(\frac{1}{n}\sum_{\ell=1}^{n}\delta_{\|\theta_{t(i)}-\theta_{t(\ell)}\|_{2}},\frac{1}{m}\sum_{\ell=1}^{m}\delta_{\|\eta_{s(j)}-\eta_{s(\ell)}\|_{2}}\right)\\ &=W_{1}\left(\sum_{k=1}^{t}\frac{n_{k}}{n}\delta_{\|\theta_{t(i)}-\theta_{k}\|_{2}},\sum_{k=1}^{s}\frac{m_{k}}{m}\delta_{\|\eta_{s(j)}-\eta_{k}\|_{2}}\right),\end{split}$
where we define
$n_{k}=\sum_{\ell=1}^{n}1\{t(\ell)=k\}\quad\text{and}\quad m_{k}=\sum_{\ell=1}^{m}1\{s(\ell)=k\}.$
We first show that $W^{\prime}(i,j)\approx\overline{W}(t(i),s(j))$ for any $(i,j)\in[n]\times[m]$. We use the fact—see, for instance, Section 2.2 of [Villani, 2003]—that the Wasserstein-1 distance on $\mathscr{P}(\mathbb{R})$ can be written as the $L^{1}$ distance between cumulative distribution functions, which yields
$\overline{W}(\alpha,\beta)$ $=\int_{\mathbb{R}}\left|\sum_{k=1}^{t}p_{k}1\{\|\theta_{\alpha}-\theta_{k}\|_{2}\leq z\}-\sum_{k=1}^{s}q_{k}1\{\|\eta_{\beta}-\eta_{k}\|_{2}\leq z\}\right|\,\mathrm{d}z\quad\forall(\alpha,\beta)\in[t]\times[s],$
$W^{\prime}(i,j)$ $=\int_{\mathbb{R}}\left|\sum_{k=1}^{t}\frac{n_{k}}{n}1\{\|\theta_{t(i)}-\theta_{k}\|_{2}\leq z\}-\sum_{k=1}^{s}\frac{m_{k}}{m}1\{\|\eta_{s(j)}-\eta_{k}\|_{2}\leq z\}\right|\,\mathrm{d}z\quad\forall(i,j)\in[n]\times[m].$
Hence,
$\begin{split}&|W^{\prime}(i,j)-\overline{W}(t(i),s(j))|\\ &\leq\int_{\mathbb{R}}\left|\sum_{k=1}^{t}\left(\frac{n_{k}}{n}-p_{k}\right)1\{\|\theta_{t(i)}-\theta_{k}\|_{2}\leq z\}\right|\,\mathrm{d}z+\int_{\mathbb{R}}\left|\sum_{k=1}^{s}\left(\frac{m_{k}}{m}-q_{k}\right)1\{\|\eta_{s(j)}-\eta_{k}\|_{2}\leq z\}\right|\,\mathrm{d}z\\ &=\int_{0}^{R}\left|\sum_{k=1}^{t}\left(\frac{n_{k}}{n}-p_{k}\right)1\{\|\theta_{t(i)}-\theta_{k}\|_{2}\leq z\}\right|\,\mathrm{d}z+\int_{0}^{R}\left|\sum_{k=1}^{s}\left(\frac{m_{k}}{m}-q_{k}\right)1\{\|\eta_{s(j)}-\eta_{k}\|_{2}\leq z\}\right|\,\mathrm{d}z\\ &\leq R\left(\sum_{k=1}^{t}\left|\frac{n_{k}}{n}-p_{k}\right|+\sum_{k=1}^{s}\left|\frac{m_{k}}{m}-q_{k}\right|\right)=:R(Z_{1}+Z_{2}).\end{split}$
Next, we show $W(i,j)\approx W^{\prime}(i,j)$ for any $(i,j)\in[n]\times[m]$. By the triangle inequality of $W_{1}$, we have
$\begin{split}&|W(i,j)-W^{\prime}(i,j)|\\ &\leq W_{1}\left(\frac{1}{n}\sum_{\ell=1}^{n}\delta_{\|X_{i}-X_{\ell}\|_{2}},\frac{1}{n}\sum_{\ell=1}^{n}\delta_{\|\theta_{t(i)}-\theta_{t(\ell)}\|_{2}}\right)+W_{1}\left(\frac{1}{m}\sum_{\ell=1}^{m}\delta_{\|Y_{j}-Y_{\ell}\|_{2}},\frac{1}{m}\sum_{\ell=1}^{m}\delta_{\|\eta_{s(j)}-\eta_{s(\ell)}\|_{2}}\right).\end{split}$
Note that we can write $X_{i}=\theta_{t(i)}+\xi_{i}$ and $Y_{j}=\eta_{s(j)}+\zeta_{j}$ for any $(i,j)\in[n]\times[m]$, where $\xi_{i}\,|\,t(i)\sim N(0,\sigma_{t(i)}^{2}I_{d})$ and $\zeta_{j}\,|\,s(j)\sim N(0,\tau_{s(j)}^{2}I_{d})$ are independent. Recall that the Wasserstein-1 distance between two sets of points in $\mathbb{R}$ with the same cardinality satisfies the following:
$W_{1}\left(\frac{1}{n}\sum_{\ell=1}^{n}\delta_{\|X_{i}-X_{\ell}\|_{2}},\frac{1}{n}\sum_{\ell=1}^{n}\delta_{\|\theta_{t(i)}-\theta_{t(\ell)}\|_{2}}\right)\leq\frac{1}{n}\sum_{\ell=1}^{n}\left|\|X_{i}-X_{\ell}\|_{2}-\|\theta_{t(i)}-\theta_{t(\ell)}\|_{2}\right|.$
Since $|\|X_{i}-X_{\ell}\|_{2}-\|\theta_{t(i)}-\theta_{t(\ell)}\|_{2}=|\|\theta_{t(i)}-\theta_{t(\ell)}+\xi_{i}-\xi_{\ell}\|_{2}-\|\theta_{t(i)}-\theta_{t(\ell)}\|_{2}|\leq\|\xi_{i}-\xi_{\ell}\|_{2}$, we have
$W_{1}\left(\frac{1}{n}\sum_{\ell=1}^{n}\delta_{\|X_{i}-X_{\ell}\|_{2}},\frac{1}{n}\sum_{\ell=1}^{n}\delta_{\|\theta_{t(i)}-\theta_{t(\ell)}\|_{2}}\right)\leq\frac{1}{n}\sum_{\ell=1}^{n}\|\xi_{i}-\xi_{\ell}\|_{2}\leq\max_{i,\ell\in[n]}\|\xi_{i}-\xi_{\ell}\|_{2}.$
Similarly,
$W_{1}\left(\frac{1}{m}\sum_{\ell=1}^{m}\delta_{\|Y_{j}-Y_{\ell}\|_{2}},\frac{1}{m}\sum_{\ell=1}^{m}\delta_{\|\eta_{s(j)}-\eta_{s(\ell)}\|_{2}}\right)\leq\max_{j,\ell\in[m]}\|\zeta_{j}-\zeta_{\ell}\|_{2}.$
Hence, we have
$|W(i,j)-W^{\prime}(i,j)|\leq\max_{i,\ell\in[n]}\|\xi_{i}-\xi_{\ell}\|_{2}+\max_{j,\ell\in[m]}\|\zeta_{j}-\zeta_{\ell}\|_{2}=:U+V.$
Fix $i\in[n]$ such that $t(i)\in[K]$. Then,
$\begin{split}\min_{\begin{subarray}{c}j\in[m]\\ s(j)\neq t(i)\end{subarray}}W(i,j)-\min_{\begin{subarray}{c}j\in[m]\\ s(j)=t(i)\end{subarray}}W(i,j)&\geq\min_{\begin{subarray}{c}j\in[m]\\ s(j)\neq t(i)\end{subarray}}W^{\prime}(i,j)-\min_{\begin{subarray}{c}j\in[m]\\ s(j)=t(i)\end{subarray}}W^{\prime}(i,j)-2(U+V)\\ &\geq\min_{\beta\in[s]\backslash\{t(i)\}}\overline{W}(t(i),\beta)-\overline{W}(t(i),t(i))-2R(Z_{1}+Z_{2})-2(U+V)\\ &\geq\Omega-2R(Z_{1}+Z_{2})-2(U+V).\end{split}$
Therefore, $2R(Z_{1}+Z_{2})+2(U+V)<\Omega$ implies that $s(\pi(i))=t(i)$ for any $i\in[n]$ such that $t(i)\in[K]$. We show that $\mathbb{P}(E)\leq\delta$, where $E$ is the event $(2R(Z_{1}+Z_{2})+2(U+V)\geq\Omega)$. By Lemma 1 and (2.3), we have
$\mathbb{P}\left(\sum_{k=1}^{t}\left|\frac{n_{k}}{n}-p_{k}\right|\geq\frac{\Omega}{8R}\right),\mathbb{P}\left(\sum_{k=1}^{s}\left|\frac{m_{k}}{m}-q_{k}\right|\geq\frac{\Omega}{8R}\right)\leq\frac{\delta}{4}.$
Since $\xi_{i}-\xi_{\ell}\,|\,t(i),t(\ell)\sim N(0,(\sigma_{t(i)}^{2}+\sigma_{t(\ell)}^{2})I_{d})$ and thus $\|\xi_{i}-\xi_{\ell}\|_{2}^{2}\,|\,t(i),t(\ell)\sim(\sigma_{t(i)}^{2}+\sigma_{t(\ell)}^{2})\chi^{2}(d)$ for $i\neq\ell$, one can deduce that
$\mathbb{P}\left(\max_{i,\ell\in[n]}\|\xi_{i}-\xi_{\ell}\|_{2}\geq\frac{\Omega}{8}\right)\leq\sum_{i\neq\ell}\mathbb{P}\left(\|\xi_{i}-\xi_{\ell}\|_{2}\geq\frac{\Omega}{8}\right)\leq n^{2}\mathbb{P}\left(\chi^{2}(d)\geq\frac{\Omega^{2}}{128\sigma^{2}}\right).$
By Lemma 2 and (2.4), we have
$\mathbb{P}\left(\max_{i,\ell\in[n]}\|\xi_{i}-\xi_{\ell}\|_{2}\geq\frac{\Omega}{8}\right)\leq n^{2}\mathbb{P}\left(\chi^{2}(d)\geq\frac{\Omega^{2}}{128\sigma^{2}}\right)\leq n^{2}\exp\left(-\frac{\Omega^{2}}{2048\sigma^{2}}\right)\leq\frac{\delta}{4}.$
Similarly, we have
$\mathbb{P}\left(\max_{j,\ell\in[m]}\|\zeta_{j}-\zeta_{\ell}\|_{2}\geq\frac{\Omega}{8}\right)\leq\frac{\delta}{4}.$
Putting all together, we have
$\mathbb{P}(E)\leq\mathbb{P}\left(Z_{1}\geq\frac{\Omega}{8R}\right)+\mathbb{P}\left(Z_{2}\geq\frac{\Omega}{8R}\right)+\mathbb{P}\left(U\geq\frac{\Omega}{8}\right)+\mathbb{P}\left(V\geq\frac{\Omega}{8}\right)\leq\delta.$
Hence, we have (2.1). ∎
Remark 1.
So far, we have focused on the matching of the points from the common part, namely, the first $K$ components. The matching guarantee ensures that all $X_{i}$’s such that $t(i)\in[K]$ are correctly matched, while ignoring the points from the outlier components. In this model, if we assume a stronger condition than (2.2), namely,
$\left(\min_{\alpha\neq\beta}\overline{W}(\alpha,\beta)\wedge\min_{\alpha\in[t]\backslash[K],\beta\in[s]}\overline{W}(\alpha,\beta)\right)-\max_{\alpha\in[K]}\overline{W}(\alpha,\alpha)=:\Omega^{\prime}>0,$ (2.5)
and replace $\Omega$ in (2.3), (2.4) with $\Omega^{\prime}$, we can show that
$\max_{i\in[n]:t(i)\in[K]}W(i,\pi(i))<\min_{i\in[n]:t(i)\notin[K]}W(i,\pi(i))$ (2.6)
holds with probability at least $1-\delta$. This means that if the threshold $\rho$ in Algorithm 1 is chosen between the two quantities in (2.6), the output $I$ of the algorithm exactly coincides with the set $\{i\in[n]:t(i)\in[K]\}$, thereby identifying the points from the common part. We provide the proof of (2.6) in Section A.1.
3 Noise Stability of Assignment-Based Matching
This section provides an additional noise stability result of the matching procedure that seeks an assignment. In the proof of Theorem 1, we have seen that $W(i,j)\approx W^{\prime}(i,j)$ which was chiefly based on the fact that the deviation of $X_{i}$ from its center $\theta_{t(i)}$ due to the noise is sufficiently small. The main principle is that the matching procedure based on distance profiles is robust to the noise added to the points. In this section, we separate this noise stability property from the previous analysis and analyze it in the context of assignment-based matching.
To this end, we consider two sets of locations $\{\theta_{1},\ldots,\theta_{n}\}$ and $\{\eta_{1},\ldots,\eta_{n}\}$ in $\mathbb{R}^{d}$ and assume $\eta_{\pi^{\ast}(i)}=T(\theta_{i})$ for $i\in[n]$, where $T\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$ is an unknown rigid transformation and $\pi^{\ast}$ is an unknown permutation on $[n]$ representing the correspondence, say, $\theta_{i}$ and $\eta_{\pi^{\ast}(i)}$ are to be matched. Unlike the previous model in Section 2, we assume there is a ground truth one-to-one correspondence $\pi^{\ast}$ between two sets of locations, which we want to recover from the following perturbed locations:
$X_{i}=\theta_{i}+\xi_{i}\quad\text{and}\quad Y_{i}=\eta_{i}+\zeta_{i},$
where $\xi_{i}\sim N(0,\sigma_{i}^{2}I_{d})$ and $\zeta_{i}\sim N(0,\tau_{i}^{2}I_{d})$ are independent noise variables. [^2] Technically, one may think of this model as the previous model in Section 2 with $t=s=K$, namely, no outliers, and with the sampling procedure that draws only one point from each component, rather than randomly drawing from $\mu,\nu$. For instance, for $d=3$, we can think of $X_{i}$’s and $Y_{i}$’s as the coordinates obtained by scanning some objects of interest but from different angles, say, frontal and lateral views, where the noise is added to the ground truth coordinates $\theta_{i}$’s and $\eta_{i}$’s during the scanning process. As the goal of this model is to recover the permutation $\pi^{\ast}$, we obtain a permutation from the following assignment-based matching procedure: defining $W(i,j)$ as in (1.1),
$\hat{\pi}=\operatorname*{arg\,min}_{\pi\in\mathcal{S}_{n}}\sum_{i=1}^{n}W(i,\pi(i)),$
which is a linear assignment problem that can be solved efficiently by polynomial-time algorithms. As discussed in Section 2, several conditions determine the success of the matching. Again, the configuration of the locations $\theta_{1},\ldots,\theta_{n},\eta_{1},\ldots,\eta_{n}$ is crucial for the matching to be well-defined. To this end, we suppose the following configuration condition:
$\Phi:=\min_{i\neq j}W_{1}\left(\frac{1}{n}\sum_{k=1}^{n}\delta_{\|\theta_{i}-\theta_{k}\|_{2}},\frac{1}{n}\sum_{k=1}^{n}\delta_{\|\theta_{j}-\theta_{k}\|_{2}}\right)>0.$ (3.1)
The condition (3.1) ensures that $\pi^{\ast}$ is the unique optimal matching in the following sense:
$\pi^{\ast}=\operatorname*{arg\,min}_{\pi\in\mathcal{S}_{n}}\sum_{i=1}^{n}W_{1}\left(\frac{1}{n}\sum_{k=1}^{n}\delta_{\|\theta_{i}-\theta_{k}\|_{2}},\frac{1}{n}\sum_{k=1}^{n}\delta_{\|\eta_{\pi(i)}-\eta_{k}\|_{2}}\right)=:\operatorname*{arg\,min}_{\pi\in\mathcal{S}_{n}}\mathcal{Q}(\pi).$
To see this, note that $\mathcal{Q}(\pi^{\ast})=0$, while $\mathcal{Q}(\pi)>0$ for $\pi\neq\pi^{\ast}$ as $\Phi>0$. The following theorem shows that the obtained permutation $\hat{\pi}$ recovers $\pi^{\ast}$ with high probability given this condition (3.1) and the variances $\sigma_{i}$’s and $\tau_{i}$’s are sufficiently small.
Theorem 2.
Suppose that the locations $\theta_{1},\ldots,\theta_{n},\eta_{1},\ldots,\eta_{n}\in\mathbb{R}^{d}$ satisfy
$\eta_{\pi^{\ast}(i)}=T(\theta_{i})\quad\forall i\in[n],$ (3.2)
for some unknown rigid transformation $T\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$ and some unknown permutation $\pi^{\ast}$ on $[n]$. Now, suppose we have the following noisy locations:
$X_{i}=\theta_{i}+\xi_{i}\quad\text{and}\quad Y_{i}=\eta_{i}+\zeta_{i},$ (3.3)
where $\xi_{i}\sim N(0,\sigma_{i}^{2}I_{d})$ and $\zeta_{i}\sim N(0,\tau_{i}^{2}I_{d})$ are independent noise variables. Let $\hat{\pi}$ be the permutation obtained from the following assignment-based matching procedure:
$\hat{\pi}=\operatorname*{arg\,min}_{\pi\in\mathcal{S}_{n}}\sum_{i=1}^{n}W_{1}(\hat{\mu}_{i},\hat{\nu}_{\pi(i)}),$
where
$\hat{\mu}_{i}:=\frac{1}{n}\sum_{\ell=1}^{n}\delta_{\|X_{i}-X_{\ell}\|_{2}}\quad\text{and}\quad\hat{\nu}_{i}:=\frac{1}{n}\sum_{\ell=1}^{n}\delta_{\|Y_{i}-Y_{\ell}\|_{2}}\quad\forall i\in[n].$
Then, for any $\delta\in(0,1)$, we have
$\mathbb{P}(\hat{\pi}=\pi^{\ast})\geq 1-\delta$
if
$\sigma^{2}\vee\tau^{2}\leq\frac{\Phi^{2}}{64(d\vee 8\log(2n^{2}/\delta))},$ (3.4)
where $\sigma=\max_{i\in[n]}\sigma_{i}$, $\tau=\max_{i\in[n]}\tau_{i}$, and $\Phi$ is defined in (3.1).
Proof of Theorem 2.
For the proof, we may assume that $\pi^{\ast}$ is the identity without loss of generality. First, letting $Q(\pi)=\sum_{i=1}^{n}W_{1}(\hat{\mu}_{i},\hat{\nu}_{\pi(i)})$, observe that
$\left(\hat{\pi}\neq\pi^{\ast}\right)\subset\bigcup_{\pi\in\mathcal{S}_{n}\backslash\{\pi^{\ast}\}}\left(Q(\pi^{\ast})\geq Q(\pi)\right).$
The event $Q(\pi^{\ast})\geq Q(\pi)$ is equivalent to
$\sum_{i=1}^{n}\left(W_{1}(\hat{\mu}_{i},\hat{\nu}_{i})-W_{1}(\hat{\mu}_{i},\hat{\nu}_{\pi(i)})\right)=\sum_{i:\pi(i)\neq i}\left(W_{1}(\hat{\mu}_{i},\hat{\nu}_{i})-W_{1}(\hat{\mu}_{i},\hat{\nu}_{\pi(i)})\right)\geq 0,$
which is contained in the following events:
$\bigcup_{i:\pi(i)\neq i}\left(W_{1}(\hat{\mu}_{i},\hat{\nu}_{i})\geq W_{1}(\hat{\mu}_{i},\hat{\nu}_{\pi(i)})\right)\subset\bigcup_{i=1}^{n}\bigcup_{j\neq i}\left(W_{1}(\hat{\mu}_{i},\hat{\nu}_{i})\geq W_{1}(\hat{\mu}_{i},\hat{\nu}_{j})\right)$
Therefore,
$\left(\hat{\pi}\neq\pi^{\ast}\right)\subset\bigcup_{i=1}^{n}\bigcup_{j\neq i}\left(W_{1}(\hat{\mu}_{i},\hat{\nu}_{i})\geq W_{1}(\hat{\mu}_{i},\hat{\nu}_{j})\right).$ (3.5)
Now, recall that we can write a rigid transformation $T\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$ as $T(\theta)=R\theta+b$ for some orthonormal matrix $R\in\mathbb{R}^{d\times d}$ and some vector $b\in\mathbb{R}^{d}$. Hence, we have
$\hat{\mu}_{i}$ $=\frac{1}{n}\sum_{\ell=1}^{n}\delta_{\|X_{i}-X_{\ell}\|_{2}}=\frac{1}{n}\sum_{\ell=1}^{n}\delta_{\|\theta_{i}-\theta_{\ell}+\xi_{i}-\xi_{\ell}\|_{2}}=\frac{1}{n}\sum_{\ell=1}^{n}\delta_{\|\theta_{i}-\theta_{\ell}+\xi_{i\ell}\|_{2}},$
$\hat{\nu}_{i}$ $=\frac{1}{n}\sum_{\ell=1}^{n}\delta_{\|Y_{i}-Y_{\ell}\|_{2}}=\frac{1}{n}\sum_{\ell=1}^{n}\delta_{\|R(\theta_{i}-\theta_{\ell})+\zeta_{i}-\zeta_{\ell}\|_{2}}=\frac{1}{n}\sum_{\ell=1}^{n}\delta_{\|\theta_{i}-\theta_{\ell}+\zeta_{i\ell}\|_{2}},$
where we let $\xi_{i\ell}=\xi_{i}-\xi_{\ell}$ and $\zeta_{i\ell}=R^{\top}(\zeta_{i}-\zeta_{\ell})$. Next, recall that the Wasserstein-1 distance between two sets of points in $\mathbb{R}$ with the same cardinality satisfies the following:
$W_{1}(\hat{\mu}_{i},\hat{\nu}_{j})=\min_{\pi\in\mathcal{S}_{n}}\frac{1}{n}\sum_{\ell=1}^{n}\left|\|\theta_{i}-\theta_{\ell}+\xi_{i\ell}\|_{2}-\|\theta_{j}-\theta_{\pi(\ell)}+\zeta_{j\pi(\ell)}\|_{2}\right|.$
Note that
$\begin{split}&\left|\left|\|\theta_{i}-\theta_{\ell}+\xi_{i\ell}\|_{2}-\|\theta_{j}-\theta_{\pi(\ell)}+\zeta_{j\pi(\ell)}\|_{2}\right|-\left|\|\theta_{i}-\theta_{\ell}\|_{2}-\|\theta_{j}-\theta_{\pi(\ell)}\|_{2}\right|\right|\\ &\leq\left|\|\theta_{i}-\theta_{\ell}+\xi_{i\ell}\|_{2}-\|\theta_{i}-\theta_{\ell}\|_{2}\right|+\left|\|\theta_{j}-\theta_{\pi(\ell)}\|_{2}-\|\theta_{j}-\theta_{\pi(\ell)}+\zeta_{j\pi(\ell)}\|_{2}\right|\\ &\leq\|\xi_{i\ell}\|_{2}+\|\zeta_{j\pi(\ell)}\|_{2}\\ &\leq(U+V),\end{split}$
where we define $U=\max_{i,\ell\in[n]}\|\xi_{i\ell}\|_{2}$ and $V=\max_{j,\ell\in[n]}\|\zeta_{j\ell}\|_{2}$. As
$W_{1}(\mu_{i},\mu_{j})=\min_{\pi\in\mathcal{S}_{n}}\frac{1}{n}\sum_{\ell=1}^{n}\left|\|\theta_{i}-\theta_{\ell}\|_{2}^{2}-\|\theta_{j}-\theta_{\pi(\ell)}\|_{2}^{2}\right|,$
we conclude that
$|W_{1}(\hat{\mu}_{i},\hat{\nu}_{j})-W_{1}(\mu_{i},\mu_{j})|\leq(U+V)\quad\forall i,j\in[n].$
Therefore, if we can find $i,j\in[n]$ such that $i\neq j$ and $W_{1}(\hat{\mu}_{i},\hat{\nu}_{i})\geq W_{1}(\hat{\mu}_{i},\hat{\nu}_{j})$, we must have
$(U+V)\geq W_{1}(\hat{\mu}_{i},\hat{\nu}_{i})\geq W_{1}(\hat{\mu}_{i},\hat{\nu}_{j})\geq W_{1}(\mu_{i},\mu_{j})-(U+V)\geq\Phi-(U+V).$
Accordingly, by (3.5), we have
$\mathbb{P}(\hat{\pi}\neq\pi^{\ast})\leq\mathbb{P}(2(U+V)\geq\Phi)\leq\mathbb{P}\left(U\geq\frac{\Phi}{4}\right)+\mathbb{P}\left(V\geq\frac{\Phi}{4}\right).$
Since $\xi_{i}-\xi_{\ell}\sim N(0,(\sigma_{i}^{2}+\sigma_{\ell}^{2})I_{d})$ and thus $\|\xi_{i\ell}\|_{2}^{2}\sim(\sigma_{i}^{2}+\sigma_{\ell}^{2})\chi^{2}(d)$ for $i\neq\ell$, one can deduce that
$\mathbb{P}\left(\max_{i,\ell\in[n]}\|\xi_{i\ell}\|_{2}\geq\frac{\Phi}{4}\right)\leq\sum_{i\neq\ell}\mathbb{P}\left(\|\xi_{i\ell}\|_{2}\geq\frac{\Phi}{4}\right)\leq n^{2}\mathbb{P}\left(\chi^{2}(d)\geq\frac{\Phi^{2}}{32\sigma^{2}}\right).$
By Lemma 2 and (3.4), we have
$\mathbb{P}\left(\max_{i,\ell\in[n]}\|\xi_{i\ell}\|_{2}\geq\frac{\Phi}{4}\right)\leq n^{2}\mathbb{P}\left(\chi^{2}(d)\geq\frac{\Phi^{2}}{32\sigma^{2}}\right)\leq n^{2}\exp\left(-\frac{\Phi^{2}}{512\sigma^{2}}\right)\leq\frac{\delta}{2}.$
Similarly, we have
$\mathbb{P}\left(\max_{j,\ell\in[n]}\|\zeta_{j\ell}\|_{2}\geq\frac{\Phi}{4}\right)\leq\frac{\delta}{2}.$
Therefore, we have $\mathbb{P}(\hat{\pi}\neq\pi^{\ast})\leq\delta$. ∎
For the case where there is no rigid transformation, namely, $T$ is the identity map, the noise stability guarantee in the above model is well-studied in the literature, known as the feature matching problem. In this case, the simplest approach is to solve the discrete optimal transport problem [Peyré and Cuturi, 2019] using the squared Euclidean distances, namely,
$\hat{\pi}^{\mathrm{OT}}=\operatorname*{arg\,min}_{\pi\in\mathcal{S}_{n}}\sum_{i=1}^{n}\|X_{i}-Y_{\pi(i)}\|_{2}^{2}.$ (3.6)
It is well-studied [Collier and Dalalyan, 2016, Galstyan et al., 2022] that such a procedure is indeed guaranteed to recover $\pi^{\ast}$ with high probability if the separation among $\theta_{i}$’s is large enough compared to the noise level $\sigma$. However, in the presence of an unknown rigid transformation $T$, such a procedure is not guaranteed to succeed, while $\hat{\pi}$ obtained in Theorem 2 is invariant to rigid transformations as we proved. We demonstrate this point in Section 4.
4 Simulations
We first provide a simulation study to demonstrate the performance of the matching procedure discussed in Section 2. Figure 2 shows the matching accuracy defined as follows:
|$\frac{|\{i\in[n]:t(i)\in[K]~{}~{}\text{and}~{}~{}s(\pi(i))=t(i)\}|}{|\{i\in[n]:t(i)\in[K]\}|},$ (4.1)
where $t,s$ are defined as in Theorem 1. In other words, we compute the proportion of correctly matched pairs among $X_{i}$’s from the first $K$ components. This accuracy measure is less strict and thus can be more practically relevant than the matching guarantee in Theorem 1 that is the probability of matching without any error. For each plot of Figure 2, we compare the case where there is one outlier, that is, $t=s=K+1$ in Assumption 1, and the baseline case where there is no outlier, namely, $t=s=K$. In the former case, $\Omega$ defined in (2.2) is $0.0345$, while $\Omega=\min_{\alpha\neq\beta}\overline{W}(\alpha,\beta)=0.1144$ for the later case as we have $\overline{W}(\alpha,\alpha)=0$ for all $\alpha\in[K]$. For both cases, we let $\sigma=\sigma_{k}=\tau_{k}$ for all $k\in[t]$ in Assumption 1 and vary $\sigma$. We can observe that the matching accuracy deteriorates as $\sigma$ grows for both cases. Though the presence of an outlier deteriorates the matching accuracy as expected, we can see that the proposed method is robust to the outlier in that the accuracy is not significantly lower than the baseline case considering the signal-to-noise ratio ($\frac{1}{K}=0.1$).
(a) $n=200$
(b) $n=500$
(c) $n=1000$
Figure 2: Accuracy of Gaussian mixture matching with $d=3$, $K=10$, and the sample size $n=m\in\{200,500,1000\}$. For each plot, the blue solid line amounts to the case where there is one outlier, that is, $t=s=K+1$ in Assumption 1, while the red dashed line is the baseline case where there is no outlier, namely, $t=s=K$. We let $\sigma_{k}=\tau_{k}=\sigma$ for all $k\in[t]$ and vary $\sigma$. For each $\sigma$, the sampling from $\mu,\nu$ is repeated 100 times and the marked points show the average of the matching accuracy over 100 realizations, while the shaded regions show the standard deviations.
Next, we provide simulation results demonstrating the noise stability discussed in Section 3. We first obtain $n$ locations $\theta_{1},\ldots,\theta_{n}\in\mathbb{R}^{d}$ and define the other locations $\eta_{1},\ldots,\eta_{n}$ as in (3.2) using some rigid transformation $T$ and a permutation $\pi^{\ast}$. Then, we generate the points according to (3.3) and obtain $\hat{\pi}$ following the procedure explained in Theorem 2; we estimate $\mathbb{P}(\hat{\pi}=\pi^{\ast})$ by repeating this for 100 times. Figure 3 shows the results, where we vary $\sigma$—assuming $\sigma=\sigma_{i}=\tau_{i}$ for all $i\in[n]$—and compare with the probability $\mathbb{P}(\hat{\pi}^{\mathrm{OT}}=\pi^{\ast})$ for $\hat{\pi}^{\mathrm{OT}}$ obtained by the procedure (3.6). As shown in Figure 3(a), when the rigid transformation is based on a rotation matrix that only affects the first two coordinates, both methods can recover the true permutation to some extent for small enough $\sigma$; in this case, as the rotation matrix is relatively close to the identity matrix, the standard optimal transport method (3.6) shows good performance. For a rotation matrix that changes all the coordinates, however, the standard optimal transport method fails to recover the true permutation as shown in Figure 3(b), while the distance profile matching method performs almost as in Figure 3(a).
(a) Changing two coordinates
(b) Changing all coordinates
Figure 3: Noise stability results with $d=10$ and $n=100$. The plots show the probability $\mathbb{P}(\hat{\pi}=\pi^{\ast})$, where $\hat{\pi}$ corresponds to the one defined in Theorem 2 for the blue lines (distance profile) and to (3.6) for the red lines (standard OT); here, the probability is estimated by repeating each setting 100 times. For (a), the rigid transformation is based on a rotation matrix that only affects the first two coordinates, while (b) is based on a random rotation matrix changing all $d$ coordinates.
5 Conclusion
In this paper, we have proposed a matching procedure based on distance profiles and established its robustness. The proposed procedure is easily implementable without having to solve nontrivial optimization problems over permutations. Most importantly, we have shown that the proposed method is provably robust to outliers and noise under some model assumptions that we believe capture essential aspects of empirical applications. To the best of our knowledge, this is the first work that provides theoretical guarantees for the robustness of a matching procedure based on distance profiles in the context of object matching. We have also demonstrated the empirical performance of the proposed method through real data and simulations.
We suggest several directions for future research. On the theoretical side, it would be interesting to consider other types of matching guarantees in Theorem 1. As mentioned in Section 4, controlling the matching accuracy (4.1) may be practically more useful. For any small $\eta\in(0,1)$, showing that the matching accuracy (4.1) is larger than $1-\eta$ with high probability would be a meaningful result. Similarly, for Theorem 2, it would be interesting to consider the probability that $\hat{\pi}$ is close to $\pi^{\ast}$ under a certain discrepancy, such as the Hamming distance, instead of the probability of a perfect recovery $\hat{\pi}=\pi^{\ast}$. On the practical side, it would be interesting to delve deeper into the procedure of finding inliers, that is, the points from the common part of two objects. In Remark 1, though we have noted that the threshold $\rho$ can be chosen to identify the inliers at least in theory, developing a practical method applicable to real data is very challenging in general. We leave these directions for future research.
Appendix A Appendix
A.1 Details of Remark 1
The proof of (2.6) in Remark 1 is similar to the proof of Theorem 1. First, recall that the event $C=(2R(Z_{1}+Z_{2})+2(U+V)<\Omega)$ implies $s(\pi(i))=t(i)$ for any $i\in[n]$ such that $t(i)\in[K]$. Hence, on the event $C$, one can deduce that
$\max_{i\in[n]:t(i)\in[K]}W(i,\pi(i))\leq\max_{\alpha\in[K]}\overline{W}(\alpha,\alpha)+R(Z_{1}+Z_{2})+(U+V).$
Meanwhile,
$\begin{split}\min_{i\in[n]:t(i)\notin[K]}W(i,\pi(i))&=\min_{i\in[n]:t(i)\notin[K]}\min_{j\in[m]}W(i,j)\\ &\geq\min_{\alpha\in[t]\backslash[K],\beta\in[s]}\overline{W}(\alpha,\beta)-R(Z_{1}+Z_{2})-(U+V).\end{split}$
Now, let $C^{\prime}:=(2R(Z_{1}+Z_{2})+2(U+V)<\Omega^{\prime})$. From (2.2) and (2.5), we have $\Omega^{\prime}\leq\Omega$ and thus $C^{\prime}\subset C$. Therefore, on the event $C^{\prime}$, we have
$\min_{i\in[n]:t(i)\notin[K]}W(i,\pi(i))-\max_{i\in[n]:t(i)\in[K]}W(i,\pi(i))\geq\Omega^{\prime}-2R(Z_{1}+Z_{2})-2(U+V)>0.$
Now, we have $\mathbb{P}(C^{\prime})\geq 1-\delta$ by repeating the proof of Theorem 1 with $\Omega$ in (2.3), (2.4) replaced by $\Omega^{\prime}$. Therefore, we have (2.6) with probability at least $1-\delta$.
A.2 Lemmas
We have used the following lemmas on the concentration of the multinomial distribution and the chi-squared distribution in the proofs of Theorems 1 and 2. The former is from [Devroye, 1983], while the latter is a standard result that can be deduced from the Chernoff bound; for instance, see Chapter 2 of [Vershynin, 2018].
Lemma 1 (Lemma 3 of [Devroye, 1983]).
Suppose $(z_{1},\ldots,z_{t})$ is from the multinomial distribution with the number of trials $n$ and the probability vector $(p_{1},\ldots,p_{t})\in\Delta_{t}$. Then, for any $\delta\in(0,1)$ such that $\delta\leq 3e^{-4t/5}$, we have
$\mathbb{P}\left(\sum_{k=1}^{t}\left|\frac{z_{k}}{n}-p_{k}\right|\geq 5\sqrt{\frac{\log(3/\delta)}{n}}\right)\leq\delta.$
Lemma 2.
Let $\chi^{2}(d)$ be the chi-squared random variable with $d$ degrees of freedom. Then,
$\mathbb{P}\left(\chi^{2}(d)\geq d+t\right)\leq\exp\left(-\frac{t(t\wedge d)}{8d}\right)\quad\forall t\geq 0.$
Accordingly, for any $t\geq 2d$,
$\mathbb{P}\left(\chi^{2}(d)\geq t\right)\leq\exp\left(-\frac{t}{16}\right).$